# LAG: LLM agents for Leaderboard Auto Generation on Demanding

### v1

@@ Abstract
This paper introduces Leaderboard Auto Generation (LAG), a novel and well-organized framework for automatic generation of leaderboards on a given research topic in rapidly evolving fields like Artificial Intelligence (AI). Faced with a large number of AI papers updated daily, it becomes difficult for researchers to track every paper’s proposed methods, experimental results, and settings, prompting the need for efficient automatic leaderboard construction. While large language models (LLMs) offer promise in automating this process, challenges such as multi-document summarization, leaderboard generation, and experiment fair comparison still remain under exploration. LAG solves these challenges through a systematic approach that involves the paper collection, experiment results extraction and integration, leaderboard generation, and quality evaluation. Our contributions include a comprehensive solution to the leaderboard construction problem, a reliable evaluation method, and experimental results showing the high quality of leaderboards.
[…]
LAG: LLM agents for Leaderboard Auto Generation on Demanding
Jian Wu1,  Jiayu Zhang2thanks: The authors contributed equally to this work.,  Dongyuan Li3,  Linyi Yang4,  Aoxiao Zhong5, Renhe Jiang3,  Qingsong Wen5,  Yue Zhang6 1 Institute of Science Tokyo 2 Peking University, China 3 University of Tokyo 4 University College London 5 AI Research Institute, Squirrel AI Learning, China 6 School of Engineering, Westlake Univeristy
@@ 1 Introduction
The explosive growth of scientific publications has created both unprecedented opportunities and significant challenges for researchers seeking to stay abreast of state-of-the-art methods Bornmann et al. (2020); Wang et al. (2024); Şahinuç et al. (2024). Leaderboard platforms, such as NLP-progress [^1] https://nlpprogress.com/ and Papers-With-Code [^2] https://paperswithcode.com/ have become invaluable by offering comprehensive overviews of recent research developments, highlighting ongoing trends, and identifying future directions. However, the large amount of daily papers makes it increasingly difficult to update these leaderboards automatically and promptly. Figure 1 illustrates two pressing issues: First, the number of LLM-related articles submitted to arXiv has surged dramatically—from 2022 to 2025, with over 20,000 submissions in 2024 alone. Second, even as new methods continuously emerge, leaderboards, such as the one for Multi-hop Question Answering on the HotpotQAYang et al. (2018) dataset, remain stagnant, with the latest method dating back to 2023. These observations highlight a serious issue: the rapid accumulation of daily scientific publications often outpaces the capability of researchers to keep up with cutting-edge research and state-of-the-art methods, emphasizing the growing need for more efficient methods to generate the latest and useful leaderboards.
@@
Figure 1: Left: Growth trend of paper submission on LLMs from 2022 to 2025-02. Right: An Example of a Multi-hop QA dataset leaderboard (HotpotQA Homepage), the latest method is still stuck in 2023.
Prior efforts have attempted to address this gap. A line of work (Hou et al., 2019; Kardas et al., 2020) has proposed leaderboard construction methods that directly extract scientific entities from individual NLP papers, and construct a static leaderboard without updating and maintenance. Semi-supervised scientific NER, proposed by Li et al. (2023), focuses on extracting scientific entities from both tables and text. Şahinuç et al. (2024) introduce SCILEAD, a manually-curated Scientific Leaderboard dataset, including 27 leaderboards derived from 43 NLP papers. However, all previous methods have been limited to the extracted scientific entities and only give a static snapshot after extracting information from a narrow selections.
We consider using LLMs such as GPT-4  Achiam et al. (2023), Qwen  Yang et al. (2024), and O1-preview which have demonstrated exceptional performance across diverse NLP tasks, especially in the long-context scenario Chen et al. (2023a, b); Wang et al. (2023b) to automatically generate leaderboards based on given research topic.
@@
Directly applying LLMs to this task still faces several key challenges. First, Limited Paper Coverage: It is challenging for human to search for all papers on a certain scientific topic, due to the overwhelming number of constantly emerging publications. Second, Unfair Comparison: Current studies do not consider fair experiment settings when making comparisons. For example, in NLP research, key experimental components, model size, train dataset size, and hyperparameter selection, vary significantly across publications, highlighting the need for automatic alignment. Finally, Low Timeliness: A leaderboard, which lacks regular updates and continuous maintenance, cannot provide researchers with sufficient useful information.
@@
To this end, we introduce LAG, a novel agent framework for dynamically and automatically generating leaderboards. Figure 2 illustrates the framework of our method, which is organized into four stages: (1) Paper Collection and split: Initially, LAG automatically download all relevant LaTex code based on the given research topic from arXiv and filter out papers published before certain date and those unrelated to the topic, ensuring proper paper coverage and timeliness. (2) Table Extraction and Classification: We use LLMs to extract and classify experiment tables based on accompanying table descriptions. (3) Table Unpacking and Integration: LAG extracts the datasets, metrics, experiment settings, and experiment results from the tables in the form of a quintuple, including paper title. Experiment setting extraction is crucial for enabling fair comparisons across different baselines. (4) Leaderboard Generation and Evaluation: The extracted quintuples are recombined and re-ranked to form candidate leaderboards.
@@
To evaluate the performance of LAG, we propose two key quality dimensions for assessment: (1) Topic-related Quality: paper coverage assessment, determining whether each quintuple in the LAG-generated leaderboards is related to the given topic. (2) Content Quality: We adopt the LLM-as-Judge method for leaderboard quality evaluation on four aspects, including Coverage, Structure, Latest, and Multiaspect. We also introduce human experts to manually evaluate LAG-generated leaderboards and compute the Pearson Correlation Coefficient between human- and LLMs-assigned scores.
Related Work Data Source Experiment Settings Multi Document Dynamic TDMS(Hou et al., 2019) NProg. Axcell (Kardas et al., 2020) PwC TELIN (Yang et al., 2022) PwC ORKG(KABENAMUALU et al., 2023) PwC LEGO (Singh et al., 2024) PwC SciLead (Şahinuç et al., 2024) NLP papers LAG (Ours) arXiv
@@
Table 1: Comparison of related work and ours. Data Source: Source of leaderboards: NProg.: NLP-progress, PwC: paperswithcode. Experiment Settings: whether the experiment settings are extracted as part of leaderboards or not. Multi Document: whether the leaderboards are constructed from multiple papers or not. Dynamic: whether the generated leaderboards can be updated dynamically or not.
@@
Extensive experiments across different leaderboard lengths (5, 10, 15, and 20 items) show that LAG consistently achieves high topic-related and content quality scores. With 20 items, a LAG-generated Leaderboard represents 20 baselines for researchers, achieving 67.58% recall and 70.33% precision scores in topic-related quality. In content quality with 20 items, LAG achieves 4.12 coverage, 3.96 latest, 4.16 structure, and 4.08 multiaspect scores, approaching human performance (4.72 coverage, 4.68 latest, 4.34 structure, and 4.58 multiaspect scores). Although the manually created leaderboard achieves higher content quality, it is much more time-consuming than LAG, deeming the efficiency. With fewer items, LAG gets even higher performance, slightly lower than human performance. These results highlight the effectiveness of LAG, providing a reliable proxy for human judgment across varying leaderboard items. Furthermore, the Pearson correlation coefficient values indicate a moderate positive correlation between the human-assigned and LLM-assigned scores.
To our best knowledge, LAG is the first method to explore the potential of LLM agents for automatic leaderboard generation, proposing evaluation criteria that align with human preferences and offering valuable reference for future related research.
@@ 2 Related Work
In the realm of LLMs, several studies have explored using LLMs for improving work efficiency in scientific research. Baek et al. (2024) and Yang et al. (2023) proposed a multi-agent-based scientific idea generation method to boost AI-related research. To evaluate the quality of LLM-generated ideas, Si et al. (2024) introduced a comprehensive human evaluation metric. Wang et al. (2023a) proposed SciMON, a method that uses LLMs for scientific literature retrieval. Wang et al. (2024) proposed an AutoSurvey to automatically generate scientific surveys based on the given research topic. The AI Scientist, Lu et al. (2024) introduced a fully automated, prompt-driven research pipeline. To make LLM-generated ideas more diverse and practical, Weng et al. (2024) proposed CycleResearcher, an iterative self-rewarding framework that allows the LLM to refine its ideas continuously, enhancing both diversity and practicality in research proposal generation. However, no previous research focused on the leaderboard generation for researchers to search, organize, and compare the state-of-the-art methods rapidly and fairly based on a certain research topic.
Figure 2: The LAG framework for leaderboard automatic generation. In Stage 1, we automatically crawl scientific papers from arXiv. In Stage 2, we retrieve, extract, and classify tables from the latex code. In Stage 3, we select the main results tables and extract datasets, metrics, results, and experiment settings from the main results table. In Stage 4, we generate Leaderboards from the selected results and evaluate the quality.
Leaderboard Construction.
Table 1 illustrates the differences between the previous work and LAG. First of all, previous work builds leaderboards by using data sources such as NLP-progress or Papers-With-Code. However, these sources lack rigorous quality assurance, such as standardizing scientific entities across different leaderboards and ensuring complete coverage of relevant publications. Instead, we choose arXiv, which is a free distribution service and an open-access archive for nearly 2.4 million scholarly articles in different domains, providing a large amount of publications for researchers. Similar to our work, Hou et al. (2019), Kardas et al. (2020), and Singh et al. (2024) extract “Task”, “Dataset”, “Model” along with the experiment result entities as TDM triples to build a leaderboard. Yang et al. (2022) and KABENAMUALU et al. (2023) leverage the pre-defined TDM triples in an extraction process similar to Hou et al. (2019). Since these approaches require a pre-defined taxonomy of TDM triples, they are incompatible with realistic task definitions. In short, none of the previous work is adaptable to constantly emerging benchmarks driven by new research and innovation. Moreover, none of the studies extract the experiment settings as additional information to generate leaderboards, which results in a lack of fair comparison. In scientific research, experiment settings are important for educators or users to reproduce the experimental results claimed in scientific publications. In this work, we address the aforementioned problems. Specifically, we (1) dynamically download scientific publications and generate up-to-date leaderboards based on the given scientific topic and the specific date; (2) extract experiment settings as part of leaderboards for fair comparison; (3) apply a Multi-Agent-as-Judge to evaluate leaderboard quality.
3 Methods
@@
Figure 2 depicts LAG, which consists of four stages: Paper Collection and Split, Table Extraction and Classification, Table Unpacking and Integration, and Leaderboard Generation and Evaluation. Each stage is meticulously designed to address specific challenges associated with leaderboard generation, thereby enhancing the efficiency and quality of the resulting leaderboards. The whole process is iterated several times (e.g., five times) to generate a high-quality leaderboard.
3.1 Paper Collection and Split
@@
Utilizing the off-the-shelf tools  [^3] https://github.com/lukasschwab/arxiv.py, LAG first searches and retrieves a set of papers $P_{\text{init}}=\{P_{1},P_{2},...,P_{N}\}$ from arXiv and downloads LaTeX code files related to a specific scientific research topic $T$. Then, we specify a certain date and filter out all papers published before the date. The filtering stage is important for ensuring that the generated leaderboards are grounded in the most relevant and recent research. Moreover, since the search tool just identifies only the keywords in the paper title and abstract, which can lead to a significant amount of noisy data, we also introduce a retrieval model to filter out papers that are irrelevant to the given topic and retrieve topic-related papers. The set of filtered papers $P_{\text{filtered}}=\{\text{Retrieval}\{P_{1},P_{2},...,P_{U}\}\}$ is used to generate the leaderboards, ensuring comprehensive coverage of the topic and logical structure. Due to the extensive number of relevant papers retrieved and filtered during this stage, the total input length of $P_{\text{filtered}}$ often exceeds the maximum input length of LLMs. Since most of the LaTex content is unproductive for generating leaderboards, we split the LaTeX code into several sections based on the structure of each paper. Most tables, table-related descriptions, experiment results, and experiment settings are located in the “Experiment” section, which contains the key information for generating leaderboards. Consequently, we select all “Experiment” sections as well as all tables $\{\text{Table}_{1},\text{Table}_{2},...,\text{Table}_{U}\}$ and all table-related descriptions $\{D_{1},D_{2},...D_{U}\}$, extracted from all papers, as input for the next stage.
3.2 Table Extraction and Classification
@@
Typically, a scientific paper, such as those in the natural language processing domain, contains several types of tables, including “Main Results”, “Ablation Study”, and “Others”. The “Main Results” tables are the most important tables in the paper, which illustrate the novelty, contributions, and effectiveness of the proposed methods or models by comparing the experiment results of the proposed method with other baselines. We utilize these tables for leaderboard generation. The “Ablation Study” tables examine the effect of damaging or removing certain components in a controlled setting to investigate all possible outcomes of system failure. The “Others” tables are the tables that illustrate the supplementary information of the experiments. For example, some tables illustrate the dataset statistics of the benchmark used in the experiments, while other tables list the results of “Case Study” and “Error Analysis”. To address this, we propose an agent that uses the In-Context Learning method Dong et al. (2022) to manually select one table from each of the three different types. The agent then prompts LLMs to classify the table types and only keeps the “Main Experiments” tables and their descriptions as the final input. The $i_{th}$ table types can be described as: $\texttt{LLM}(\text{Table}_{i},D_{i};\text{Prompt})\rightarrow\texttt{Table type}$.
@@
In practice, the most intrinsic approach is to divide Stage 2 into the following sequential steps: (1) Extract all tables and their associated captions from the LaTeX code. (2) Classify the extracted tables according to predefined table types. (3) Extract metrics, performance values, and experimental settings related to the proposed model from tables categorized as “Main Results”. Moreover, each of these three steps necessitates the use of LLM APIs, and repeated reference to certain table contents further exacerbates the substantial waste of tokens. To address this issue, we create the agent following the few-shot Chain of Thought (CoT) prompting process, enabling it to classify and extract information from identified “Main Results” tables in a single dialogue round. Specifically, in the requested JSON output, we additionally set the key points as follows: “number of tables (Int)", “classification of tables (Dict)" and “selected table’s index (Int)".
3.3 Table Unpacking and Integration

### v2

# League: Leaderboard Generation on Demand

@@ Abstract
Leaderboard gathers experimental results from various sources into unified rankings, giving researchers clear standards for measuring progress while facilitating fair comparisons. However, with thousands of academic papers updated daily, manually tracking each paper’s methods, results, and experimental settings has become impractical, creating an urgent need for automated leaderboard generation. Although large language models offer promise in automating this process, challenges such as multi-document summarization, fair result extraction, and consistent experimental comparison remain underexplored. To address these challenges, we introduce Leaderboard Auto Generation (League), a novel and well-organized framework for automatic generation of leaderboards on a given research topic in rapidly evolving fields like Artificial Intelligence. League employs a systematic pipeline encompassing paper collection, result extraction and integration, leaderboard generation, and quality evaluation. Through extensive experiments across multiple research domains, we demonstrate that League produces leaderboards comparable to manual curation while significantly reducing human effort.  [^1] Code will be available upon publication.
@@ 1 Introduction
The explosive growth of scientific publications has created both unprecedented opportunities and significant challenges for researchers seeking to stay abreast of state-of-the-art methods (Bornmann et al., 2020; Wang et al., 2024; Şahinuç et al., 2024). Leaderboard platforms, such as NLP-progress [^2] https://nlpprogress.com/ and Papers-With-Code [^3] https://paperswithcode.com/ have become invaluable by offering comprehensive overviews of recent research developments, highlighting ongoing trends, and identifying future directions. However, the overwhelming growth of daily papers makes it increasingly difficult to build and update leaderboards automatically and promptly. Figure 1 illustrates two pressing issues: 1) a widening gap: LLM-related paper submissions to arXiv have surged year-over-year (exceeding 55,000 in 2025), yet the growth trend of leaderboard submissions on Paper with Code (5670 in 2025) has not advanced accordingly. 2) Even as new methods continuously emerge, leaderboards, such as the one for Multi-hop Question Answering on the HotpotQA(Yang et al., 2018) dataset, remain stagnant, with the latest method dating back to 2023. These observations highlight a serious challenge: the rapid accumulation of daily scientific publications often outpaces the capability of researchers to keep up with cutting-edge research and state-of-the-art methods, emphasizing the growing need for more efficient methods to generate the latest and useful leaderboards.
@@
Figure 1: Left: Growth trend of paper and leaderboard submissions on LLMs from 2022 to 2025-09. The leaderboard statistics are collected from Paper with Code. Right: Leaderboard of Multi-hop QA, the latest method is still stuck in 2023.
Earlier studies have made initial attempts to tackle this challenge. Research on scientific information extraction (Hou et al., 2019; Kardas et al., 2020) has focused on identifying entities such as models, datasets, metrics, and results from individual NLP papers, which enables the creation of paper-specific leaderboards. However, such approaches are inherently difficult to maintain and update over time. More recently, Li et al. (2023), proposed Scientific NER to extract entities from both text and tables, while Şahinuç et al. (2024) introduced SCILEAD, a hand-curated dataset containing 27 leaderboards from 43 NLP papers for evaluating LLMs on entity extraction. Despite these advances, prior methods remain restricted to entity extraction, offering only preliminary building blocks for leaderboard creation and yielding static snapshots from a narrow set of papers. In contrast, our work shifts the focus to dynamic leaderboard construction, where the goal is not only to extract results but also to continuously track, align, and compare model performance on specific datasets or tasks under standardized evaluation settings. This enables researchers to observe evolving performance trends and conduct fair comparisons, thereby providing a more comprehensive and up-to-date view of progress in a research area.
@@
Directly applying LLMs such as GPT-4 (Achiam et al., 2023), Qwen (Yang et al., 2024), and O1-preview that have demonstrated exceptional performance in various NLP tasks, especially in the long-context scenario (Chen et al., 2023a; b; Wang et al., 2023b) to this task faces several key challenges. First, Limited Paper Coverage: It is challenging for humans to search for all papers on a certain scientific topic, due to the overwhelming number of constantly emerging publications. Second, Unfair Comparison: Current studies do not consider fair experiment settings when making comparisons. For example, in NLP research, key experimental components, model size, train dataset size, and hyperparameter selection vary significantly across publications, highlighting the need for automatic alignment. Finally, Low Timeliness: A leaderboard, which lacks regular updates and continuous maintenance, cannot provide researchers with sufficient useful information.
@@
To address these issues, we introduce League, a novel framework for dynamically and automatically leaderboard generation. Figure 2 illustrates the framework of our method, which is organized into four stages: (1) Paper Collection and split: Initially, League automatically downloads all relevant LaTex and PDF files based on the given research topic from arXiv and top-tier conferences, including ACL, EMNLP, NeurIPS, ICML, and ICLR, and filters out papers published before a certain date and those unrelated to the topic, ensuring proper paper coverage and timeliness. (2) Table Extraction and Classification: We use LLMs to extract and classify experiment tables based on accompanying table descriptions. (3) Table Unpacking and Integration: League extracts the datasets, metrics, experiment settings, and experiment results from the tables in the form of a quintuple, including paper titles. Experiment setting extraction is crucial for enabling fair comparisons across different baselines (including model size, data size, etc). (4) Leaderboard Generation and Evaluation: The extracted quintuples are recombined, refined, and re-ranked to form candidate leaderboards.
@@
We assess performance along two dimensions: (1) Topic-related Quality: whether each quintuple in League-generated leaderboards relates to the given topic; (2) Content Quality: evaluated via LLM-as-Judge on four aspects including Coverage, Structure, Latest, and Multiaspect. Human experts also evaluate the leaderboards, with Pearson Correlation computed between human and LLM scores.
@@
Table 1: Comparison between related work and our approach. Data Source: Source of leaderboards. NProg.: NLP-progress, PwC: paperswithcode. Experiment Settings: whether the experiment settings are extracted as part of leaderboards. Multi Document: whether the leaderboards are constructed from multiple papers. Dynamic: whether the generated leaderboards can be updated dynamically.
Related Work Data Source Exp. Settings Multi Doc. Dynamic Timeliness TDMS(Hou et al., 2019) N Prog. Axcell (Kardas et al., 2020) PwC TELIN (Yang et al., 2022) PwC ORKG(KABENAMUALU et al., 2023) PwC LEGO (Singh et al., 2024) PwC SciLead (Şahinuç et al., 2024) NLP papers League (Ours) arXiv
@@
Experiments across different leaderboard lengths (5, 10, 15, and 20 items) show that League consistently achieves high Topic-related and Content Quality scores, approaching human performance but about 5-10 times the speed of human annotation on 20-item leaderboard construction. With 20 items, a League-generated Leaderboard represents 20 baselines for researchers, achieving 67.58% recall and 70.33% precision scores in Topic-related Quality. In Content Quality with 20 items, League achieves 4.12 coverage, 3.96 latest, 4.16 structure, and 4.08 multiaspect scores, approaching human performance (4.72 coverage, 4.68 latest, 4.34 structure, and 4.58 multiaspect scores). Although the manually created leaderboard achieves higher Content Quality, it is much more time-consuming than League, highlighting League’s superior efficiency. With fewer items, League gets even higher performance, slightly lower than human performance. These results highlight the effectiveness of League, providing a reliable proxy for human judgment across varying leaderboard items. Furthermore, the Pearson Correlation Coefficient values indicate a moderate positive correlation between the human-assigned and LLM-assigned scores. To the best of our knowledge, we are the first to explore the potential of LLMs for automatic leaderboard generation, proposing evaluation criteria that align with human preferences and offering valuable reference for future related research.
@@ 2 Related Work
Several studies have explored using LLMs to improve work efficiency in scientific research  (Xie et al., 2025). Baek et al. (2024) and Yang et al. (2023) propose a multi-agent-based scientific idea generation method to boost AI-related research. To evaluate the quality of LLM-generated ideas, Si et al. (2024) introduces a comprehensive human evaluation metric. Wang et al. (2023a) proposes SciMON, a method that uses LLMs for retrieving the scientific literature. Wang et al. (2024) proposes an AutoSurvey to automatically generate scientific surveys based on the given research topic. The AI Scientist (Lu et al., 2024) introduces a fully automated and prompt-driven research pipeline. To make LLM-generated ideas more diverse and practical, Weng et al. (2024) proposes an iterative self-rewarding framework that allows the LLM to continuously refine its ideas, improving both diversity and practicality in research proposal generation. However, no previous research focused on leaderboard generation for researchers to search, organize, and compare the state-of-the-art methods rapidly and fairly based on a certain research topic.
Scientific Information Extraction.
Table 1 illustrates the differences between the previous experiment results extraction work and League. Earlier works on scientific information extraction mainly focused on extracting entities such as Task, Dataset, and Model (TDM triples) from sources like NLP-progress or Papers-with-Code (Hou et al., 2019; Kardas et al., 2020; Singh et al., 2024). Others (Yang et al., 2022; KABENAMUALU et al., 2023) extended this approach by leveraging predefined TDM taxonomies. However, these methods face three key limitations: (1) they rely on incomplete or inconsistently curated sources, leading to gaps in coverage; (2) they depend on fixed taxonomies, making them inflexible to new benchmarks or emerging research directions; and (3) they only extract entities without capturing experimental settings, preventing fair and reproducible comparisons. As a result, prior efforts can at best provide static snapshots of research progress, but fail to maintain up-to-date, comparable leaderboards.
In contrast, our League framework goes beyond mere extraction: it automatically crawls arXiv and top conference proceedings, identifies and unpacks main experimental tables, extracts both results and experiment settings, and then integrates these quintuples to produce complete, up‑to‑date leaderboards. By combining dynamic paper collection, table classification, settings-aware result integration, and LLM‑based refinement and evaluation, League generates leaderboards end‑to‑end rather than only extracting isolated result entities. This shift enables timely, reproducible, and fair comparisons across methods and makes League suitable for rapidly evolving research areas where static extraction methods fall short.
3 Method
@@
Figure 2 depicts the proposed League, which consists of four stages: (1) Paper Collection and Split, (2) Table Extraction and Classification, (3) Table Unpacking and Integration, and (4) Leaderboard Generation and Evaluation. Each stage is meticulously designed to address specific challenges associated with leaderboard generation, thereby enhancing the efficiency and quality of the resulting leaderboards. The whole process is iterated several times (e.g., five times) to generate a high-quality leaderboard.
Figure 2: The League framework for leaderboard automatic generation. In Stage 1, we automatically crawl scientific papers from arXiv. In Stage 2, we retrieve, extract, and classify tables from the LaTeX code. In Stage 3, we select the main results tables and extract datasets, metrics, results, and experiment settings from the main results table. In Stage 4, we generate Leaderboards from the selected results and evaluate the quality.
3.1 Stage I Paper Collection and Split
@@
Utilizing the off-the-shelf tools [^4] https://github.com/lukasschwab/arxiv.py, League first searches and retrieves a set of papers $P_{\text{init}}=\{P_{1},P_{2},...,P_{N}\}$ from arXiv and top-tier conferences and downloads LaTeX or PDF files related to a specific research topic $T$. Then, we specify a certain date and filter out all papers published before the date. The filtering stage is important for ensuring that the generated leaderboards are grounded in the most relevant and recent research. Moreover, since the search tool just identifies only the keywords in the paper title and abstract, which can lead to a significant amount of noisy data, we also introduce a retrieval model to filter out papers that are irrelevant to the given topic and retrieve topic-related papers. The set of filtered papers $P_{\text{filtered}}=\{\text{Retrieval}\{P_{1},P_{2},...,P_{U}\}\}$ is used to generate the leaderboards, ensuring comprehensive coverage of the topic and logical structure. Due to the extensive number of relevant papers retrieved and filtered during this stage, the total input length of $P_{\text{filtered}}$ often exceeds the maximum input length of LLMs. Since most of the LaTeX content is unproductive for generating leaderboards, we split the LaTeX code into several sections based on the structure of each paper. Most tables, table-related descriptions, experiment results and experiment settings are located in the “Experiment” section, which contains the key information to generate leaderboards. Thus, we select all “Experiment” sections as well as all tables $\{\text{Table}_{1},\text{Table}_{2},...,\text{Table}_{U}\}$ and all table-related descriptions $\{D_{1},D_{2},...D_{U}\}$, extracted from all papers, as input for the next stage.
3.2 Stage II Table Extraction and Classification
@@
Typically, a scientific paper, such as those in the natural language processing domain, contains several types of tables, including “Main Results”, “Ablation Study”, and “Others”. The “Main Results” tables are the most important tables in the paper, which illustrate the novelty, contributions, and effectiveness of the proposed methods or models by comparing the experiment results of the proposed method with other baselines. We utilize these tables for leaderboard generation. The “Ablation Study” tables examine the effect of damaging or removing certain components in a controlled setting to investigate all possible outcomes of system failure. The “Others” tables are the tables that illustrate the supplementary information of the experiments. For example, some tables illustrate the dataset statistics of the benchmark used in the experiments, while other tables list the results of the “Case Study” and “Error Analysis”. To address this, we propose a framework that uses the In-Context Learning method (Dong et al., 2022) to manually select one table from each of the three different types. The framework then prompts LLMs to classify the table types and only keeps the “Main Experiments” tables and their descriptions as the final input. The $i_{th}$ table types can be described as: $\texttt{LLM}(\text{Table}_{i},D_{i};\text{Prompt})\rightarrow\texttt{Table type}$.
@@
In practice, the most intrinsic approach is to divide Stage 2 into the following sequential steps: (1) Extract all tables and their associated captions from the LaTeX code. (2) Classify the extracted tables according to predefined table types. (3) Extract metrics, performance values, and experimental settings related to the proposed model from tables categorized as “Main Results”. However, each of these three steps necessitates the use of LLM APIs, and repeated reference to certain table contents further exacerbates the substantial waste of tokens. To address this issue, we follow the few-shot Chain of Thought (CoT) prompting process, enabling it to classify and extract information from identified “Main Results” tables in a single dialogue round. Specifically, in the requested JSON output, we additionally set the key points as follows: “number of tables (Int)", “classification of tables (Dict)", and “selected table’s index (Int)".
3.3 Stage III Table Unpacking and Integration